OpenStudy (anonymous):

Solve by completing the square and then use the square root method to solve. Write your answer in set brackets with no spaces, such as {-1,2}. x^2 - 4x - 32 = 0

OpenStudy (anonymous):

\[x^2-4x=32\] \[(x-2)^2=32+2^2=32+4=36\] \[x-2=\pm\sqrt{36}\] \[x-2=\pm6\] \[x-2=6\iff x=8\] \[x-2=-6\iff x=-4\]

OpenStudy (anonymous):

How would i put that in brackets ?